OpenStudy (anonymous):

find the solution. dy/dt = (ty)^2

OpenStudy (anonymous):

SO you want to integrate?

OpenStudy (anonymous):

well separate variables and integrate solve for y

OpenStudy (anonymous):

Do you have the answers on an answer sheet or something?

OpenStudy (anonymous):

I just need check if I got it right.

OpenStudy (anonymous):

need to*

OpenStudy (anonymous):

yea y(t) = -3/(t^3 + k)

OpenStudy (anonymous):

So where is the problem, you need the steps?

OpenStudy (anonymous):

Don't misunderstand my question please, but I think you mentioned everything that's necessary already above, this is an ODE, it is separable.

OpenStudy (anonymous):

i just need the steps. i dont know how you get that answer

OpenStudy (whpalmer4):

\[dy/dt = (ty)^2\]\[y^{-2}*dy = t^2dt\]\[\int y^{-2}dy = \int t^2 dt\]\[-y^{-1} + c_1 = \frac{1}{3}t^3 + c_2\]\[y^{-1} = -\frac{1}{3}t^3 + c\]Take reciprocal of both sides \[y = - \frac{1}{\frac{1}{3}t^3 + c}\]\[y = -\frac{3}{t^3+c}\]

OpenStudy (anonymous):

I don't get your last line @whpalmer4 How can you get a three on the numerator with a 1/3t^3 +c in the denominator?

OpenStudy (anonymous):

(1/3)

OpenStudy (whpalmer4):

Ah, I just factor out the 1/3 and turn it into 3 in the numerator. The values of c obviously aren't identical, but one constant is as good as another, I always say :-)

OpenStudy (anonymous):

Yeah. so you can just make 3c into c, correct? Cause multiplying the constant woulds till give you a constant but the constant would be much larger?

OpenStudy (whpalmer4):

Right. What I'm really doing is \[y = -\frac{1}{\frac{1}{3}t^3 + c} = -\frac{1}{\frac{1}{3}(t^3 + 3c)} = -\frac{3}{t^3+3c} = -\frac{3}{t^3+k}\]

OpenStudy (whpalmer4):

\[k = 3c\]

OpenStudy (anonymous):

thank you whpalmer4. that was very helpful

OpenStudy (whpalmer4):

Glad I could still remember how to do it! It's been a while :-)

OpenStudy (whpalmer4):

As a little exercise, take the derivative of the result and plug it back into the original to make sure it works (it does, but it's good practice).